(c17p18) Use the following values for mass and charge: an electron has mass me = 9.11
Answer #1

since

KE=1/2mv2=qV=1.6x10-19 x 750=1.2x10-16J

since

1/2mv2=qV

v=sqrt(2qV/m)=sqrt (2 x1.6x10-19 x 750/9.11x10-31)=1.623x107m/s